# Scalarisation-based risk concepts for robust multi-objective optimisation

### v1

@@ Abstract
Robust optimisation is a well-established framework for optimising functions in the presence of uncertainty. The inherent goal of this problem is to identify a collection of inputs whose outputs are both desirable for the decision maker, whilst also being robust to the underlying uncertainties in the problem. In this work, we study the multi-objective extension of this problem from a computational standpoint. We identify that the majority of all robust multi-objective algorithms rely on two key operations: robustification and scalarisation. Robustification refers to the strategy that is used to marginalise over the uncertainty in the problem. Whilst scalarisation refers to the procedure that is used to encode the relative importance of each objective. As these operations are not necessarily commutative, the order that they are performed in has an impact on the resulting solutions that are identified and the final decisions that are made. This work aims to give an exposition on the philosophical differences between these two operations and highlight when one should opt for one ordering over the other. As part of our analysis, we showcase how many existing risk concepts can be easily integrated into the specification and solution of a robust multi-objective optimisation problem. Besides this, we also demonstrate how one can principally define the notion of a robust Pareto front and a robust performance metric based on our robustify and scalarise methodology. To illustrate the efficacy of these new ideas, we present two insightful numerical case studies which are based on real-world data sets.
@@ 1 Introduction
Robust multi-objective optimisation is concerned with the problem of optimising a vector-valued function under uncertainty. Precisely, we consider optimising a bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, which is defined over a $D$-dimensional input space $\mathbb{X}\subseteq\mathbb{R}^{D}$ and a $W$-dimensional uncertain parameter space $\Xi\subseteq\mathbb{R}^{W}$. Here, the inputs $\mathbf{x}\in\mathbb{X}$ are used to denote all the parameters that can be controlled by the decision maker. Whilst the uncertain parameters $\boldsymbol{\xi}\in\Xi$ are used to denote all the other variables in the problem, which effect the value of the objective function, but are subject to uncertainty. Notably, these uncertain parameters might not be controllable. As a result, the goal of robust multi-objective optimisation is to identify a collection of controllable inputs whose distribution of output vectors are in some sense robust to the potential variation in these uncertain parameters. That is, we aim to solve the problem
@@
where the ‘robustmax’ operator is some sensible robust generalisation of the vector-valued maximum which has to be defined by the decision maker. The majority of existing work in robust multi-objective optimisation has largely focussed on defining the robust optimum by taking a first principles approach where one defines a partial ordering over the space of inputs—see for instance the work by Ide and Köbis (2014); Ide and Schöbel (2016) and Botte and Schöbel (2019) for a review. In particular, in this first principles approach, one comes up with a way to say whether an input $\mathbf{x}\in\mathbb{X}$ is more robust than another input $\mathbf{x}^{\prime}\in\mathbb{X}$ with respect to the objective function $f$. The solution to the corresponding robust optimisation problem is then given by the collection of inputs which are not robustly dominated by any other feasible input. Consequently, after a notion of robustness has been established, one is typically interested in coming up with a principled computational strategy in order to identify these robust points. Overall, the most prevalent approach to accomplish this is the scalarisation strategy. In this approach, one recasts the corresponding robust multi-objective optimisation problem into a sensible collection of scalar-valued optimisation problems that can be solved using traditional techniques. More specifically, one typically tries to identify a collection of scalar-valued optimisation problem whose solutions are necessarily robust in the desired sense. Although theoretically motivated, this first principles approach is rather rigid in practice because whenever we want to use a new partial ordering, we would have to manually identify and derive a new set of results. In this work, we flip the script and take a more computationally-driven approach to robust multi-objective optimisation. Instead of defining robustness from first principles, we define robustness directly using scalarisation functions. That is, we define the solution of the robust multi-objective optimisation (1) as being equivalent to solving a collection of robust scalarised problems. By construction, this class of robust problems is very attractive because it can be generalised very easily and can be readily solved using standard scalar-valued optimisation techniques.
@@
There are the two essential ingredients to our robust multi-objective optimisation strategy: robustification and scalarisation. Robustification is concerned with the process of quantifying the robust value of a set of possible points. Whilst scalarisation is concerned with the computational process of actually solving for the best possible input which achieves the best robust value. In this work, we focus on the use of risk functionals, otherwise known as risk measures, in order to perform the robustification procedure. These risk inspired concepts of robustness are very well-established in the robust single-objective optimisation literature (Ben-Tal et al., 2009; Bertsimas et al., 2011; Rahimian and Mehrotra, 2019), but have not been adequately studied in the multi-objective setting. Our work aims to address this gap by providing a general methodology where one can easily slot in whichever risk functional best suits their problem. On the other side, for the scalarisation procedure, we opt to keep the discussion rather general and do not impose any constraints on the scalarisation mappings. Nevertheless, as a working example, we focus on the length scalarisation functions, which are a variant of the Chebyshev scalarisation function that is known to exhibit many intuitive and useful properties (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024b). As identified in preceding papers such as the works by Fliege and Werner (2014) and Groetzner and Werner (2022), the robustification and scalarisation operations are not commutative in general: the order matters. A core part of our work is a careful analysis of when one should robustify then scalarise (RTS) or whether one should scalarise then robustify (STR). Philosophically, as we highlight later in Section 3.4.1, one should take an RTS approach in the many-time aggregating scenario where one is interested in the aggregated performance at an input over many repeated runs. In contrast, one should take an STR approach in the one-time or non-aggregating scenario where one wishes to obtain consistently strong performance on every individual run.
@@ 1.1 Main contributions
As part of this work, we have made several novel and insightful contributions to the study of robust multi-objective optimisation. Firstly, we have introduced and formalised two general computationally-driven approaches to robust multi-objective optimisation: the RTS approach (Section 3.2) and the STR approach (Section 3.3). As we motivate in Section 3.1, these two approaches are very flexible and encompasses many existing robust multi-objective solution strategies as special cases. Secondly, we present a useful discussion on when one should use an RTS or STR approach (Section 3.4.1) and when these approaches are equivalent (Section 3.4.2). Thirdly, inspired by the recent results by Tu et al. (2024b), we showcase in Section 4 how it is possible to define a valid Pareto front surface which reflects the robust trade-off that is happening under both the RTS and STR paradigm for the length scalarised problems. Specifically, we derive the novel concept of the RTS front (Definition 4.3) and STR front (Definition 4.4), which are polar surfaces described by the robust length scalarised values. Fourthly, in Section 5, we present a general class of performance metrics that can be used under the RTS and STR setting in order to assess the performance of any robust multi-objective algorithm. In particular, we generalise the construction of the R2 utility (Hansen and Jaszkiewicz, 1998) to the RTS and STR setting. Lastly, in Section 6, we demonstrate the usefulness of all of our new ideas on some practical numerical experiments which are based on real-world data sets.
@@ 1.2 Structure of the paper
The rest of the paper is organised as follows: In Section 2, we recall the basics of multi-objective optimisation (Section 2.1), the scalarisation strategy (Section 2.2) and risk functionals (Section 2.3). We then move on to Section 3, where we introduce the robust multi-objective optimisation problem and recall the key ideas from existing work (Section 3.1). Afterwards, we present the RTS approach (Section 3.2) and STR approach (Section 3.3) and then include a discussion on when these approaches are appropriate and when they are equivalent (Section 3.4). We then move on to Section 4, where we introduce two generalisations of the Pareto front surface, namely the RTS front and STR front. Besides this, in Section 5, we introduce the robust R2 utilities, which can be used to quantify the performance of any approximate robust Pareto set. Consequently, in Section 6, we present some numerical experiments that showcase some of the keys ideas of our methodology in practice. Finally, in Section 7, we conclude the paper with a short summary and discussion. All of the proofs for the main results in this paper are presented in Appendix A, whilst the scripts that are needed to reproduce the figures and experiments are available in our Github repository: https://github.com/benmltu/scalarize.
@@ 2 Preliminaries
This section provides a focussed recap on the basics of multi-objective optimisation (Section 2.1), the scalarisation approach (Section 2.2) and risk functionals (Section 2.3).
@@ 2.1 Multi-objective optimisation
where the maximum here is defined using the Pareto partial ordering relations, which we recall below. Note that in this work, we focus entirely on maximisation problems. This can be done without loss of generality because any minimisation problem can be turned into a maximisation problem via a standard negation argument: $\min(A)=-\max(-A)$ for any set of vectors $A\subset\mathbb{R}^{M}$.
@@
The Pareto domination region is defined as the collection of points which dominates (or is dominated) by a particular set of vectors $A\subset\mathbb{R}^{M}$, that is
@@ 2.2 Scalarisation perspective
The standard strategy that is often used to solve the multi-objective optimisation problem (2) is the scalarisation method. In this approach, we recast the vector-valued optimisation problem of interest into a collection of scalar-valued optimisation problems
$\max_{\mathbf{x}\in\mathbb{X}}s_{\boldsymbol{\theta}}(g(\mathbf{x})),$ (3)
where $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a parametric scalarisation function that is indexed by scalarisation parameters $\boldsymbol{\theta}\in\Theta$. Notably, if we had a scalarised optimisation problem (3) for every Pareto optimal point, then solving this latter collection of problems is in some sense equivalent to solving the former. As a result, this class of problems is much more general and flexible than the original multi-objective problem (2). In particular, one can formulate scalarised problems where the target solutions are desirable to the decision maker but not necessary Pareto optimal. This feature is very useful in a preference-driven scenario where the goal of interest is to identify points that achieve a certain specification. For instance, one might use a collection of distance-based scalarisation function (Example 2.1) in order to target a collection of desirable points.
@@
If a decision maker could elicit a desired target vector $\boldsymbol{\upsilon}\in\mathbb{R}^{M}$, then they might consider setting the scalarisation function to be the negative weighted $L^{p}$ distance to this point:
@@
for some weight vector $\mathbf{w}\in\Delta^{M-1}:=\{\mathbf{y}\in\mathbb{R}_{\geq 0}^{M}:||\mathbf{y}||_{L^{1}}=1\}$. The negative sign is included because we want the points closer to the target vectors to have a higher scalarised value. Note that the solution to the corresponding scalarised optimisation problem (3) is not necessarily Pareto optimal if the target vector lies in the feasible space. Specifically, the scalarisation function (4) would not be monotonic over the feasible space and it would actually penalise objective vectors $g(\mathbf{x})\in\mathbb{R}^{M}$ which Pareto dominate the target vector $\boldsymbol{\upsilon}$.
Monotonicity
There are several results that describe the properties that a scalarisation function must satisfy for the solutions to be Pareto optimal—see for instance Miettinen (1998) or Ehrgott (2005) for a review. An often quoted desirable property is monotonicity (Tu et al., 2024a, Definition 2.5). Notably, whenever a scalarisation function is monotonically increasing, then it is known that there exists some solutions to the corresponding scalarised problem (3) which are Pareto optimal in some sense. This property is beneficial in the setting where the goal of interest is indeed to target the Pareto optimal points. Although, as pointed out above, there might be scenarios where the decision maker’s preferences are not monotonic and therefore the additional flexibility offered by the scalarisation approach over the standard Pareto ordered approach (2) is very attractive.
@@
Weak Pareto front
In Theorem 2.1, we recall a key result from the multi-objective optimisation literature, which states that all weakly Pareto optimal solutions of the multi-objective problem (2) can be recovered by solving a set of Chebyshev scalarised problems (3). Precisely, we focus our attention on a specific reparameterisation of the Chebyshev scalarisation function, which has been referred to as the achievement or length scalarisation function (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024b):
@@
for any vector $\mathbf{y}\in\mathbb{R}^{M}$, where $\boldsymbol{\eta}\in\mathbb{R}^{M}$ denotes a fixed reference vector and $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$ denotes a positive unit vector that lives in the space of all positive unit vectors $\mathcal{S}^{M-1}_{+}:=\{\mathbf{y}\in\mathbb{R}_{>0}^{M}:||\mathbf{y}||_{L^{2}}=1\}$. Conceptually, for any vector $\mathbf{y}$ that strongly dominates the reference vector $\boldsymbol{\eta}$, the length scalarisation function returns the projected length of this vector along the line $\{\boldsymbol{\eta}+t\boldsymbol{\lambda}:t\in\mathbb{R}\}$, that is, $s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(\mathbf{y})=\max\{t>0:\boldsymbol{\eta}+t\boldsymbol{\lambda}\preceq\mathbf{y}\}$ for any $\mathbf{y}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$—see the work by Tu et al. (2024b) for more details for this construction. For some additional intuition, we have included Figure 1, which gives a pictorial depiction of how the length scalarisation function can be used to target any weakly Pareto optimal point in two dimensions.
@@
Remark 2.1 (Strict solutions)
In some cases, one might be more interested in targetting just the strictly Pareto optimal solutions, which is clearly a subset of the weakly Pareto optimal solutions: $\mathcal{X}_{g}^{\textnormal{strict}}\subseteq\mathcal{X}_{g}^{\textnormal{weak}}$. To accomplish this, one can introduce an additional penalty term into the length scalarised problem (6). Informally, this additional penalty term works by penalising any points which lie on the ‘flat’ parts of the weak Pareto front. These points naturally correspond to the output of any improper Pareto optimal solutions. That is, any solution whose output behaves like a point that is weakly Pareto optimal, but not strictly Pareto optimal. More details on these penalised adaptation are described in the work by Miettinen (1998, Part 2, Section 3.4.5).

### v2

@@ Abstract
Robust optimisation is a well-established framework for optimising functions in the presence of uncertainty. The inherent goal of this problem is to identify a collection of inputs whose outputs are both desirable for the decision maker, whilst also being robust to the underlying uncertainties in the problem. In this work, we study the multi-objective case of this problem. We identify that the majority of all robust multi-objective algorithms rely on two key operations: robustification and scalarisation. Robustification refers to the strategy that is used to account for the uncertainty in the problem. Scalarisation refers to the procedure that is used to encode the relative importance of each objective to a scalar-valued reward. As these operations are not necessarily commutative, the order that they are performed in has an impact on the resulting solutions that are identified and the final decisions that are made. The purpose of this work is to give a thorough exposition on the effects of these different orderings and in particular highlight when one should opt for one ordering over the other. As part of our analysis, we showcase how many existing risk concepts can be integrated into the specification and solution of a robust multi-objective optimisation problem. Besides this, we also demonstrate how one can principally define the notion of a robust Pareto front and a robust performance metric based on our “robustify and scalarise” methodology. To illustrate the efficacy of these new ideas, we present two insightful case studies which are based on real-world data sets.
@@ 1 Introduction
Robust multi-objective optimisation is concerned with the problem of optimising a vector-valued function under uncertainty. In this paper, we will consider optimising a bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, which is defined over a $D$-dimensional input space $\mathbb{X}\subseteq\mathbb{R}^{D}$ and a $W$-dimensional uncertain parameter space $\Xi\subseteq\mathbb{R}^{W}$. Here, the inputs $\mathbf{x}\in\mathbb{X}$ are used to denote all the parameters that can be controlled by the decision maker. Whilst the uncertain parameters $\boldsymbol{\xi}\in\Xi$ are used to denote all the other variables in the problem, which effect the value of the objective function, but are subject to uncertainty or randomness. Notably, these uncertain parameters might not be controllable. As a result, the goal of robust multi-objective optimisation is to identify a collection of controllable inputs whose distribution of output vectors are in some sense robust to the potential variation in these uncertain parameters. That is, we aim to solve the problem
@@
where the ‘robustmax’ operator is some sensible robust generalisation of the vector-valued maximum which has to be defined by the decision maker.
@@
The majority of existing work in robust multi-objective optimisation has largely focussed on defining the robust optimum via a first principles Pareto ordering style approach where one defines a partial ordering over the space of inputs—see for instance the work by Ide and Köbis (2014); Ide and Schöbel (2016) and Botte and Schöbel (2019) for a review. Conceptually, this partial ordering determines whether an input $\mathbf{x}\in\mathbb{X}$ is considered to be more robust than another input $\mathbf{x}^{\prime}\in\mathbb{X}$ with respect to the objective function $f$. The solution to the corresponding robust optimisation problem is then given by the collection of inputs which are not robustly dominated by any other feasible input. Consequently, one is then typically interested in coming up with a principled computational strategy in order to identify these robust points. The most prevalent approach to accomplish this is the scalarisation strategy. In this approach, one recasts the corresponding robust multi-objective optimisation problem into a sensible collection of scalar-valued optimisation problems which can then be solved using traditional techniques in order to obtain an approximation to the robust set of interest. Although theoretically motivated, the first principles approach is rather rigid in practice because whenever we want to use a new partial ordering, we would have to manually identify and derive a new set of results. In this work, we take a more efficient approach to robust multi-objective optimisation and define robustness directly using scalarisation functions. That is, we define the solution of the robust multi-objective optimisation (1) as being equivalent to solving a collection of general robust scalarised problems that can be readily solved using standard scalar-valued optimisation techniques.
@@
There are the two essential ingredients to our robust multi-objective optimisation strategy: scalarisation and robustification. Scalarisation refers to the procedure of mapping a vector-valued objective function to a family of scalar-valued rewards. Whilst robustification is concerned with the process of quantifying the robust value of a set of possible points. In this work, we focus on the use of risk functionals, otherwise known as risk measures, in order to perform the robustification procedure. These risk inspired concepts of robustness are very well-established in the robust single-objective optimisation literature (Ben-Tal et al., 2009; Bertsimas et al., 2011; Rahimian and Mehrotra, 2019), but have not been adequately studied in the multi-objective setting. Our work aims to address this gap by providing a generic methodology for general risk functionals. Similarly, for the scalarisation procedure, our methodology is general and can accommodate for any arbitrary family of scalarisation functions. Nevertheless, as a working example, we will often use the length scalarisation functions. These are a variant of the Chebyshev scalarisation functions that are known to exhibit many intuitive and useful properties which will be discussed throughout (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024b).
As identified in preceding papers (Fliege and Werner, 2014; Groetzner and Werner, 2022), the robustification and scalarisation operations are not commutative in general: the order matters. A core part of our work is a careful analysis of when one should robustify then scalarise (RTS) or whether one should scalarise then robustify (STR). We will highlight later in Section 3.3.2, that one should take an RTS approach in the many-time aggregating scenario where one is interested in the aggregated performance at an input over many repeated runs. In contrast, one should take an STR approach in the one-time or non-aggregating scenario where one wishes to obtain consistently strong performance on every individual run.
@@ 1.1 Main contributions
As part of this work, we have made several novel and insightful contributions to the study of robust multi-objective optimisation. Firstly, we have introduced and formalised two general computationally efficient approaches to robust multi-objective optimisation: the RTS approach (Section 3.1) and the STR approach (Section 3.2). As we motivate in Section 3.3.1, these two approaches are very flexible and they encompass many existing robust multi-objective solution strategies as special cases. Secondly, we present a detailed discussion on when one should use an RTS or STR approach (Section 3.3.2) and when these approaches are equivalent (Section 3.3.3). Thirdly, inspired by the recent results by Tu et al. (2024b), we showcase in Section 4 how it is possible to define a valid Pareto front surface which reflects the robust trade-off that is happening under both the RTS and STR paradigm for the length scalarised problems. Specifically, we derive the novel concept of the RTS front (Definition 4.3) and STR front (Definition 4.4), which are polar surfaces described by the robust length scalarised values. Fourthly, in Section 5, we present a general class of performance metrics that can be used under the RTS and STR setting in order to assess the performance of any robust multi-objective algorithm. In particular, we generalise the construction of the R2 utility (Hansen and Jaszkiewicz, 1998) to the RTS and STR setting.
@@ 1.2 Structure of the paper
The rest of the paper is organised as follows: In Section 2, we recall the basics of multi-objective optimisation (Section 2.1), the scalarisation strategy (Section 2.2) and risk functionals (Section 2.3). We then move on to Section 3, where we first introduce the RTS approach (Section 3.1) and STR approach (Section 3.2) and then discuss their relation with existing work (Section 3.3). Afterwards, we move on to Section 4, where we introduce two generalisations of the Pareto front surface, namely the RTS front and STR front. Besides this, in Section 5, we introduce the robust R2 utilities, which can be used to quantify the performance of any approximate robust Pareto set. Consequently, in Section 6, we present some numerical experiments which showcases how our methodology can be used in practice on some real-world data sets. Finally, in Section 7, we conclude the paper with a short summary and discussion. All of the proofs for the main results in this paper are presented in Appendix A, whilst the scripts that are needed to reproduce the figures and experiments are available in our Github repository: https://github.com/benmltu/scalarize.
@@ 2.1 Multi-objective optimisation
where the maximum here is defined using the Pareto partial ordering relations, which we recall below. Note that throughout, we will use $g(\mathbf{x})\in\mathbb{R}^{M}$ to denote a generic vector-valued objective; we will use $f(\mathbf{x},\boldsymbol{\xi})\in\mathbb{R}^{M}$ when the objective is also dependent on an uncertainty parameter $\boldsymbol{\xi}\in\Xi$. In addition, we will focus entirely on maximisation problems. This can be done without loss of generality as $\min(A)=-\max(-A)$ for any set of vectors $A\subset\mathbb{R}^{M}$.
@@
The Pareto domination region of a set of vectors $A\subset\mathbb{R}^{M}$ is defined as the collection of vectors which dominates (or is dominated by) this set, that is
@@ 2.2 Scalarisation perspective
The standard strategy that is often used to solve (2) is the scalarisation method. In this approach, we recast the vector-valued optimisation problem of interest into a collection of scalar-valued optimisation problems
$\{\max_{\mathbf{x}\in\mathbb{X}}s_{\boldsymbol{\theta}}(g(\mathbf{x})):\boldsymbol{\theta}\in\Theta\},$ (3)
where $s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a parametric scalarisation function that is indexed by scalarisation parameters $\boldsymbol{\theta}\in\Theta$. Notably, if one is interested in capturing the Pareto front using (3), then this would require us to define a suitable set of scalarisation functions $\{s_{\boldsymbol{\theta}}\}_{\boldsymbol{\theta}\in\Theta}$ which ensures that the resulting solutions are Pareto optimal. For instance, this property is usually assured by restricting our attention to scalarisation functions that are monotonically increasing (Tu et al., 2024a, Definition 2.5). In practice however, a decision maker might not necessarily be interested in just targetting the Pareto front. In particular they might be open to the use of non-monotonic scalarisation functions, which can be used to target desirable points that are not necessarily Pareto optimal. For example, they might consider the use of distance-based scalarisation functions in order to target a set of desired specifications.
@@
If a decision maker could elicit a desired target vector $\boldsymbol{\upsilon}\in\mathbb{R}^{M}$, then they might consider setting the scalarisation function to be the negative [^1] The negative sign is included because we want the points closer to the target vectors to have a higher scalarised value. weighted $L^{p}$ distance to this point:
@@
for some weight vector $\mathbf{w}\in\Delta^{M-1}:=\{\mathbf{y}\in\mathbb{R}_{\geq 0}^{M}:||\mathbf{y}||_{L^{1}}=1\}$. Note that the solution to the corresponding scalarised optimisation problem (3) is not necessarily Pareto optimal if the target vector lies in the feasible space. Specifically, the scalarisation function (4) would not be monotonic over the feasible space and it would actually penalise objective vectors $g(\mathbf{x})\in\mathbb{R}^{M}$ which Pareto dominate the target vector $\boldsymbol{\upsilon}$.
@@
As a working example, we will often use the length scalarisation function, which is a specific reparameterisation of the Chebyshev scalarisation function (Ishibuchi et al., 2009; Deb and Jain, 2012; Tu et al., 2024b):
@@
for any vector $\mathbf{y}\in\mathbb{R}^{M}$, where $\boldsymbol{\eta}\in\mathbb{R}^{M}$ denotes a fixed reference vector and $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$ denotes a positive unit vector that lives in the space of all positive unit vectors $\mathcal{S}^{M-1}_{+}:=\{\mathbf{y}\in\mathbb{R}_{>0}^{M}:||\mathbf{y}||_{L^{2}}=1\}$. Conceptually, for any vector $\mathbf{y}$ that strongly dominates the reference vector $\boldsymbol{\eta}$, the length scalarisation function returns the projected length of this vector along the line $\{\boldsymbol{\eta}+t\boldsymbol{\lambda}:t\in\mathbb{R}\}$, that is, $s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(\mathbf{y})=\max\{t>0:\boldsymbol{\eta}+t\boldsymbol{\lambda}\preceq\mathbf{y}\}$ for any $\mathbf{y}\in\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$—see the work by Tu et al. (2024b) for more details for this construction. We now recall a key result from the multi-objective optimisation literature, which states that all weakly Pareto optimal solutions of (2) can be recovered by solving a set of length scalarised problems (3). For some additional intuition, we have also included Figure 1, which showcases the geometric idea behind this result in two dimensions.